OpenStudy (anonymous):

Please help! I'll fan you! (square root of 4-6)^2=16. why?

OpenStudy (rahulmr):

\[\sqrt{4-6}=-4\] and \[-4^2\]=16. Hope this make sense

OpenStudy (anonymous):

It does! Thanks!!

563blackghost (563blackghost):

Nice @Rahulmr

563blackghost (563blackghost):

@meranda123 you may medal people who helped you greatly....just click on "Best response" to show your thanx toward @Rahulmr Also since your question has been answered you would close your question so there wont be a mess on the open question.....to close it you would click on "Close" right at the bottom of where your question is presented....Alsooooo \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) \(\Large{By:thomaster}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct